find the y intercept and x intercept for the equation
2x-4y=16
order of pairs please

Answer by bucky(2189)

Given the equation 2x - 4y = 16
.
You are asked to find the y intercept and the x intercept.
.
It helps to understand that any coordinate point on the y axis has 0 as the value of x. Therefore, to find the value of the y intercept, just go to the equation and set x equal to zero. When you do that, the equation becomes:
.
(2*0) - 4y = 16
.
and this reduces to:
.
-4y = 16
.
Solve for y by dividing both sides by -4 and you get:
.
y = 16/-4 = -4
.
So the y intercept is the point (0, -4)
.
In a similar way, any coordinate point on the x axis has zero as its value for y. Therefore, to find the x intercept, go to the equation and set y equal to zero. When you do that the equation is:
.
2x - (4*0) = 16
.
And this reduces to:
.
2x = 16
.
Solve for x by dividing both sides by 2 and you have:
.
x = 16/2 = 8
.
So the x intercept is the coordinate point (8, 0)
.
Hope this helps you to understand graphing and intercepts a little better.